OpenStudy (anonymous):

How are e and ln related? If I have e^(4x) and x=ln2, where can I go from there? (This is calc related)

OpenStudy (jamesj):

f(x) = e^x and g(x) = ln(x) are inverse functions of each other: \[ f(g(x)) = e^{\ln x} = x\] \[ g(f(x)) = \ln(e^x) = x \]

OpenStudy (jamesj):

Therefore, if \( x = \ln 2 \), then \[ 4x = 4 \ln 2 = \ln(2^4) = \ln(16) \] and hence \[ e^{4x} = ... what? \]

OpenStudy (anonymous):

It's equal to 16, then?

OpenStudy (jamesj):

Yes

OpenStudy (anonymous):

Awesome! Thanks!